OpenStudy (anonymous):

HELP! i give medals! Simplify!! x^4-4x^2/x^2-9 ÷ 4x^2-4x^3+x^4/x^2-6x+9

OpenStudy (anonymous):

\[\frac{x^4-4x^2}{x^2-9} \div \frac{4x^2-4x^3+x^4}{x^2-6x+9}\]

OpenStudy (anonymous):

All polynomials will need to be factored, then change the division to multiplication and invert the second fraction. Cancel all common factors and multiply.

OpenStudy (anonymous):

\[\frac{x^2(x+2)(x-2)}{(x+3)(x-3)} \div \frac{4x^2-4x^3+x^4}{x^2-6x+9}\]